OpenStudy (anonymous):

What is the least common denominator of 5/12 and 11/18?? @DatChinookGuy A. 6 B. 36 C. 72 D. 144

OpenStudy (anonymous):

The given denominators are 8 and 16 The multiples of 8: 8, 16, 24, 32, 40 and so on... The multiples of 16: 16, 32, 48 and so on... Therefore you may notice for both of the denominators the least common number is 16. The least common denominator of 18 and 116 are 16. The given denominators are 4 and 9 The multiples of 4: 4, 8, 12, 16, 20, 24, 28, 32, 36, 38 and so on... The multiples of 9: 9, 18, 27, 36, 45 and so on... you may notice for both of the denominators the least common number is 36. the least common denominator of 14 and 19 is 36.

OpenStudy (anonymous):

So the answer would be B 36

OpenStudy (anonymous):

Thanks a lot @lerodge857

OpenStudy (anonymous):

Anytime :D I work for medals though :P jk

OpenStudy (amorfide):

it is also the first number the 12 and 18 times table share

OpenStudy (anonymous):

@amorfide I do not know the way u did it!

OpenStudy (amorfide):

i think my way only worked coz they are both in the 6 times table probs better to delete it

OpenStudy (anonymous):

o okay lol

OpenStudy (anonymous):

@dewdroplyss UGHH sorry I was AFK and totally missed that you tagged me. anyways, ================================================ 12 = 2 x 2 x 3 18 = 2 x 3 x 3 7 = 7 20 = 2 x 2 x 5 LCD = 2 x 2 x 3 x 3 x 5 x 7 = 1260 Convert all fractions to have an LCD of 1260: 5 x ( 1260 / 12) / 1260 = 525 / 1260 11 x (1260 / 18) / 1260 = 770 / 1260 4 x (1260 / 7) / 1260 = 720 / 1260 9 x (1260 / 20) / 1260 = 567 / 1260 So from smallest to largest: 525/1260 (5 / 12) 567/1260 (9/20) 720/1260 (4/7) 770/1260 (11/18) x ()() Take the prime factor of 12,18,7,20 and count how many times they occur in each and multiply them. --------------------------------------... Example if you don't understand: Find the LCD of 9/12, 14/18, 66/100 Prime factors of 12 are: 2, 2, and 3 Prime factors of 18 are: 2, 3, and 3 Prime factors of 100 are: 2, 2, 5, and 5 LCD = 2 x 2 x 3 x 3 x 5 x 5 = 900 Then to Find a common denominator of 12,18, 7, and 20. then however many times the bottom number goes into the common denomminator multiply that by your top number. then you have all the fractions with the same denomintor and differnt numerators so you can correctly order them. but before you put them in that new order use the defualt numbers (5/12,11/18,4/7,9/20) Therefore you may notice for both of the denominators the least common number is 16. The least common denominator of 18 and 116 are 16. The multiples of 4: 4, 8, 12, 16, 20, 24, 28, 32, 36, 38 and so on... The multiples of 9: 9, 18, 27, 36, 45 and so on... you may notice for both of the denominators the least common number is 36. the least common denominator of 14 and 19 is 36. ============================== So in the conclusion your answer would be B. 36. Hope this helps, have a nice day. SO SORRY I MISSED THAT YOU TAGGED ME! Let me know if you need help on anymore.